OpenStudy (anonymous):

find the value of (3+3i)^4

OpenStudy (anonymous):

-324

OpenStudy (bluemiku):

-324

OpenStudy (bluemiku):

thats what i got to^^

OpenStudy (shadowlegendx):

@turqoisetuna @BlueMiku It is against OpenStudy rules to give plain answers.

OpenStudy (shadowlegendx):

Please show @dabeezy how you got your answers

OpenStudy (bluemiku):

ok

OpenStudy (perl):

( 3 + i ) ^ 4 = (3 ( 1 + i ) ) ^4 = 3^4 ( 1 + i ) ^ 4 , now change 1 + i to polar trig form

OpenStudy (anonymous):

. i need a little guidance how to get there

OpenStudy (bluemiku):

oh thats what i did to

OpenStudy (midhun.madhu1987):